1For each of the following, determine the expected cash flows. (Round answers to 0 decimal places, e.g. 5,275. Show amounts th2Pharoah Inc. owns and operates a number of hardware stores in the Atlantic region. Recently, the company has decided to open
Answer #1

1

Total expected value in case of future cash flows adjusted with probability estimates are calculated as follows:

Total Expected Value = P1 * C1 + P2 * C2 + P3 * C3, wher P = Probability and C = Cash Flow

So, the answers are as follows:

a) Total Expected Value = 4000*0.3 + 6800*0.6 + 9000*0.1 = 1200 + 4080 + 900 = 6180 $

b) Total Expected Value = 5000*0.3 + 8200*0.4 + 9500*0.3 = 1500 + 3280 + 2850 = 7630$

c) Total Expected Value = (1200)*0.15 + 3500*0.65 + 5000*0.2 = (180) + 2375 + 1000 = 3195$

2 For this question the factor tables are required , else the calculations will be difficult to explain